OpenStudy (anonymous):

Consider Integral of (x^2+x)dx from 1 to 3 using Riemann sums. If u divide the interva from 1 to 3 into N subintervals, what will bethe length of each subinterval? What will be those subintervals?

OpenStudy (anonymous):

The length of each subinterval will be (total length)/(number of subintervals) Total length = top limit - bottom limit = 3 - 1 = 2 number of subintervals = N length of each subinterval = 2/N The first subinterval goes from 1 to 1 + 2/N The 2nd from 1 + 2/N to 1 + 4/N The 3rd from 1 + 4/N to 1 + 6/N The kth from 1 + 2(k-1)/N to 1 + 2k/N The nth (last) from 1 + 2(N -1)/N to 1 + 2N/N which is 3.